7) A disk is initially spinning about its center at 23 rad/s counter-clockwise and a constant angular acceleration of 2.8 rad/s2 is applied to it clockwise. How many revolutions does it make before it stops?
Answer #1

here,

initial angular speed, wi = 23 rad/s

final angular speed = 0 rad/s

angular acceleration = - 2.8 rad/s^2

let the angle is theta

Using third equation of motion

wf^2 - wi^2 = 2 * a * theta

0 - 23^2 = 2 * (-2.8) * theta

theta = 94.5 radian = 15.04 rev

the number of revolutions is 15.04 rev